OpenStudy (anonymous):

LAST QUESTION PLEASE HELP Amanda tells you that when variables are in the denominator, the equation becomes unsolvable. "There is a value for x that makes the denominator zero, and you can't divide by zero," Amanda explains. Using complete sentences, demonstrate to Amanda how the equation is still solvable. please help!!!

OpenStudy (wolf1728):

Let's suppose the equation is 1/(x-7) = 21 When x = 7 the equation is unsolvable because of the zero denominator. However, the equation is still solvable because x does not equal 7. 1/(x-7) = 21 1 = 21x -147 21x = 148 x = 148/21 x = 7.0476190476

OpenStudy (anonymous):

okay

OpenStudy (wolf1728):

Glad to help you out.